OpenStudy (anonymous):

Help me Solve Definite Integrals of the following detail: int ((3 * x ^ 3-2 * x ^ 2 +4) / ... x = 1 .. 4)?

zepdrix (zepdrix):

\[\Large\sf \int\limits \frac{3x^3-2x^2+4}{...x}dx=1..4\quad ?\]What..? What are all the dots supposed to be?

OpenStudy (anonymous):

= 39/2 ?

zepdrix (zepdrix):

??

OpenStudy (anonymous):

The result is = 39/2?

zepdrix (zepdrix):

Your question is not clear. What is the integral you're trying to solve? Why did you put a bunch of dots everywhere?

zepdrix (zepdrix):

Is this the problem? \[\Large\sf \int\limits_1^4 \frac{3x^3-2x^2+4}{x}dx\]

OpenStudy (anonymous):

missed by dividing the x ^ 2

OpenStudy (anonymous):

\[\int\limits_{1}^{4}\frac{ 3x^3-2x^2+4 }{ x^2}\]